OpenStudy (swissgirl):

What is the difference? For every positive real number x there is a positive real number y with the property that if y

OpenStudy (anonymous):

but it's not true.....

OpenStudy (swissgirl):

The first option is true and the second one is false

OpenStudy (swissgirl):

So there must be some difference that i am missing

OpenStudy (kinggeorge):

The first can be written \[\forall x\in\mathbb{R}^+ \;\;\exists y\in\mathbb{R}^+ \;\;s.t. \;\;y<x\implies (\forall z\in\mathbb{R}^+ \;\; yz \geq z)\]The second is\[\forall x\in\mathbb{R}^+ \;\;\exists y\in\mathbb{R}^+ \;\;s.t. \;\;y<x \;\text{and}\; (\forall z\in\mathbb{R}^+ \;\; yz \geq z)\]The difference is in that implication.

OpenStudy (swissgirl):

so how wld that difference make one true and one false?

OpenStudy (swissgirl):

maybe if i use it

OpenStudy (swissgirl):

its contrapositive?

OpenStudy (kinggeorge):

I'm not quite sure why that's different at first glance. Give me a couple more minutes to think about this.

OpenStudy (swissgirl):

okkk

OpenStudy (kinggeorge):

I believe the key is that y=1 is the only option for y.

OpenStudy (kinggeorge):

In the first way of writing it, there are two cases Case 1: x>1. Here, we choose y=1. We have that \(y<x\), and \(yz\geq z\)So that's taken care of. Case 2: \(x\leq1\). Here, we choose \(y\geq1\). Since this makes \(y<x\) false, it doesn't matter if \(yz\geq z\) or not, since the implication will always be true.

OpenStudy (swissgirl):

YAYYYYYYYYY

OpenStudy (kinggeorge):

In the second way of writing it, we have the same cases. The only difference is in case 2. Case 2: \(x\leq 1\). Here, we are forced to choose \(y<1\), and then we can choose a \(z\) such that \(yz>z\) which makes the statement false.

OpenStudy (swissgirl):

yup I forgot abt that one

OpenStudy (swissgirl):

THANNNKSSSSS :))))))))))

OpenStudy (kinggeorge):

You're welcome.